OpenStudy (butterflyhope):

Convert 1.52 into fraction form and reduce to lowest terms. I know this should be really easy but I'm tired and my brain is stuck ; - ; Thanks!

OpenStudy (anonymous):

38/25

OpenStudy (texaschic101):

the number to the left of the decimal is your whole number. So your whole number is 1. Now look at your .52...the 2 is in the 100ths place...so that makes your fraction 52/100. So you now have 1 52/100...turned to fraction = 152/100 reduced = 38/25